Use the Laplace transform to solve the following initial value problem:


y′′+4y=8δ(t−7)y(0)=0,y′(0)=0


y(t)=
Answer #1

Given y+4y - 88(4-7) 900)=0 gosso Apply Laplace foarsform on both sides 4[4). +4 ([y] = 8L{$(-2)] →619 [y] = ${y(s) - $460Formula il és Fis)] = ult-a) f(t-a) -- y(t) = 4 ult-7) sin (2(t-7)) = 44(4-7) (sin (24-14)]